OpenStudy (anonymous):

I am not able to prove the second part of the question , i proved the first part

OpenStudy (anonymous):

If the roots of the equation \[\huge \frac{ 1 }{ (x+p) } + \frac{ 1 }{ (x+q) } = \frac{ 1 }{ r }\] are equal in Magnitude but opposite in sign , show that :- (a)p+q = 2r and the (b) product of the roots is equal to (-1/2)(p^2 + q^2)

OpenStudy (anonymous):

@Kainui

OpenStudy (tkhunny):

How did you get the first part? If you get that, the second should rather be staring at you. Did you manage \(x^{2} + (p+q-2r)x + (qp-rp-rq) = 0\) as an equivalent equation?

OpenStudy (anonymous):

\[\huge x ^{2} + (p+q-2r)x -[\frac{ \left( p+q)r- pq \right) }{ r }]\] Let (p+q)r - pq be k (a constant) x^2 + (p+q-2r)x - k = 0 If one root is a and the other b a+b=0<------- Henceforth on substituting a and b in the equation and by quadratic formula finding out a and b then a+b = -p - q +2r 0 = -p -q +2r p+q=2r

OpenStudy (anonymous):

QED

OpenStudy (tkhunny):

Since the solutions are equal and opposite, this must represent a Difference of Squares and has to look like this. \(x^{2} - a^{2} = 0\) Thus, \(p + q - 2r = 0\;and\;p+q = 2r\) Thus, ... Hold on. How did you get 'r' in the denominator in that constant term?

OpenStudy (anonymous):

I simmplified the main equaation , ( i used wolfram alpha) and yes the constant is diided and also multiplied by r so it kindoff cancels out

OpenStudy (tkhunny):

Kinda, but not. You should have only \(-(r(p+q) - pq)\). Something went wrong in there. There shoudl be no 'r' in the denominator. That denominator should be unity (1).

OpenStudy (anonymous):

Yes that's what it is 1 after i simplify it further i missed out r earlier

OpenStudy (tkhunny):

If you agree the constant term is -(r(p+q) - pq), then we're almost done.

OpenStudy (anonymous):

It is -[(p+q)r -pq]) that 's what i got oh yes u got the same

OpenStudy (tkhunny):

From the result in Part 1, we have \(r = \dfrac{p+q}{2}\). Substitute and simplify. See what happens.

OpenStudy (anonymous):

what to substitute ? and how does that make Product of roots

OpenStudy (anonymous):

@tkhunny r u there

OpenStudy (anonymous):

@tkhunny @tkhunny @tkhunny @tkhunny @tkhunny @tkhunny R U there!

OpenStudy (tkhunny):

Are you forgetting your algebra? Roots: a and b. Factors: (x-a)(x-b) Equation: \(x^{2} -(a+b)x + ab = 0\) See how the coefficient on the x-term is the sum of the roots and the constant term is the product of the roots? That is what we need to see. You have r(p+q) + pq and you have r = (p+q)/2. Substitute the second into he first and simplify things.

OpenStudy (tkhunny):

...and, never do that again. :-(